OpenStudy (anonymous):

an object is 420 pounds when it is at sea level. Find the value of C that makes the equation true. (Sea level is 3,963 miles from the center of the Earth.) can someone help me work this out please I dont understand any of it

OpenStudy (anonymous):

w=Cr^-2

OpenStudy (anonymous):

w=420(3963)^2 w=420(15705369) w=6596254980

OpenStudy (mertsj):

You did it right.

OpenStudy (anonymous):

ok I got that part do you mind if i give you the next step and see if its right?

OpenStudy (anonymous):

so is that my c value?

OpenStudy (mertsj):

Oh. Sorry. I didn't notice that you messed up your variables. First rewrite the equation so that it says: w = C/r^2 Now multiply both sides by r^2 so you have: w(r^2) =C Then replace w with 420 lbs and r with 3963 miles, square r and multiply by 420 and you will have C.

OpenStudy (anonymous):

so will it be 420(3963)^-2=6596254980?

OpenStudy (anonymous):

no wait thats the same thing isnt it lol

OpenStudy (mertsj):

So it will be (420)(3963)^2 = C (420)(15705369)=C 6596254980 = C

OpenStudy (anonymous):

ok then my next step is to use the value of c to find how much the object would weigh is it is 450ft about sea level

OpenStudy (anonymous):

w=6596254980(3963-450/5280)^-2